# Croppable Knowledge Graph Embedding

### v1

@@ Appendix A More Results of link prediction
Method MRR Hit@1 MRR Hit@1 MRR Hit@1 MRR Hit@1 MRR Hit@1 MRR Hit@1 MRR Hit@1
TransE DT .150 .102 .277 .190 .299 .212 .313 .218 .315 .222 .318 .224 .322 .228
Ext .115 .065 .191 .122 .236 .156 .266 .180 .286 .197 .299 .208 .322 .228
Ext-L .109 .065 .175 .115 .232 .157 .263 .180 .285 .198 .301 .210 .322 .228
Ext-V .139 .081 .200 .126 .237 .156 .270 .185 .293 .205 .308 .217 .322 .228
BKD .176 .106 .279 .198 .303 .208 .315 .222 .315 .223 .320 .226 - -
TA .175 .112 .281 .200 .303 .212 .314 .220 .319 .225 .321 .223 - -
DualDE .179 .115 .281 .201 .306 .216 .316 .223 .319 .226 .322 .227 - -
MED .196 .122 .290 .199 .308 .218 .317 .223 .320 .226 .321 .227 .322 .227
RotatE DT .254 .168 .297 .207 .312 .223 .317 .224 .322 .229 .323 .230 .325 .234
Ext .138 .080 .203 .129 .251 .170 .276 .190 .291 .203 .305 .217 .325 .234
Ext-L .135 .078 .188 .121 .221 .146 .246 .166 .280 .193 .299 .209 .325 .234
Ext-V .160 .097 .198 .126 .238 .159 .265 .182 .288 .201 .302 .213 .325 .234
BKD .277 .193 .305 .214 .314 .224 .321 .230 .322 .230 .323 .231 - -
TA .280 .196 .306 .216 .313 .225 .319 .229 .323 .229 .323 .231 - -
DualDE .282 .197 .307 .216 .315 .227 .318 .230 .322 .232 .324 .233 - -
MED .288 .201 .311 .216 .318 .225 .322 .231 .323 .233 .324 .233 .324 .232
PairRE DT .182 .116 .243 .162 .284 .202 .307 .222 .319 .227 .328 .235 .332 .237
Ext .148 .107 .177 .118 .217 .149 .259 .182 .294 .207 .321 .230 .332 .237
Ext-L .150 .099 .196 .134 .219 .159 .271 .188 .309 .219 .326 .233 .332 .237
Ext-V .176 .116 .192 .125 .229 .154 .279 .193 .311 .221 .329 .237 .332 .237
BKD .198 .132 .251 .168 .288 .203 .311 .224 .321 .233 .330 .236 - -
TA .208 .139 .263 .182 .292 .210 .314 .224 .323 .232 .332 .235 - -
DualDE .207 .139 .261 .179 .293 .212 .316 .226 .326 .234 .334 .238 - -
MED .239 .172 .274 .189 .303 .213 .314 .224 .324 .232 .329 .236 .330 .235
Table 11: Hit@10 and Hit@3 of some representative dimensions on FB15K237.
@@
Method Hit@10 Hit@3 Hit@10 Hit@3 Hit@10 Hit@3 Hit@10 Hit@3 Hit@10 Hit@3 Hit@10 Hit@3 Hit@10 Hit@3
TransE DT .235 .169 .440 .301 .477 .327 .484 .340 .499 .348 .501 .353 .508 .358
Ext .211 .123 .324 .211 .392 .264 .436 .296 .462 .320 .479 .331 .508 .358
Ext-L .194 .118 .293 .192 .381 .256 .424 .292 .462 .316 .484 .333 .508 .358
Ext-V .256 .150 .348 .222 .396 .265 .437 .301 .466 .325 .488 .341 .508 .358
BKD .293 .178 .446 .308 .480 .336 .500 .349 .501 .349 .502 .354 - -
TA .246 .188 .441 .307 .484 .336 .498 .348 .504 .353 .504 .355 - -
DualDE .301 .193 .443 .307 .483 .337 .502 .351 .505 .354 .508 .356 - -
MED .341 .215 .472 .321 .486 .338 .502 .347 .505 .351 .507 .356 .507 .358
RotatE DT .424 .284 .477 .330 .495 .346 .502 .352 .506 .353 .510 .357 .515 .363
Ext .245 .152 .340 .225 .410 .278 .443 .304 .465 .322 .485 .335 .515 .363
Ext-L .243 .147 .319 .209 .365 .247 .402 .275 .453 .312 .477 .333 .515 .363
Ext-V .281 .174 .340 .218 .393 .264 .427 .293 .458 .319 .478 .336 .515 .363
BKD .442 .306 .485 .338 .503 .352 .508 .354 .510 .356 .509 .358 - -
TA .447 .308 .485 .339 .501 .353 .507 .358 .510 .359 .509 .358 - -
DualDE .449 .311 .486 .341 .502 .353 .507 .360 .512 .361 .514 .361 - -
MED .459 .324 .492 .344 .504 .355 .509 .357 .510 .358 .512 .362 .514 .362
PairRE DT .314 .198 .395 .262 .452 .312 .476 .337 .505 .352 .518 .364 .522 .368
Ext .222 .158 .289 .187 .353 .236 .416 .283 .469 .325 .506 .354 .522 .368
Ext-L .249 .159 .294 .196 .333 .238 .436 .298 .489 .342 .513 .359 .522 .368
Ext-V .277 .181 .303 .192 .374 .250 .450 .307 .490 .343 .513 .362 .522 .368
BKD .332 .215 .407 .265 .453 .314 .487 .343 .508 .355 .521 .366 - -
TA .346 .226 .430 .291 .455 .316 .493 .347 .509 .358 .521 .368 - -
DualDE .342 .224 .427 .286 .456 .318 .495 .351 .512 .359 .524 .371 - -
MED .384 .253 .437 .299 .466 .327 .495 .346 .510 .357 .521 .366 .520 .368
@@
(b) RotatE on WN18RR
(c) PairRE on WN18RR
(d) TransE on FB15K237
(e) RotatE on FB15K237
(f) PairRE on FB15K237
@@ B.2 Evolutionary Improvement Mechanism (EIM)
$L_{KGE}^{i}=\sum_{(h,r,t)\in\mathcal{T}\cup\mathcal{T}^{-}}y\log\sigma(s_{(h,r,t)}^{i})+(1-y)\log(1-\sigma(s_{(h,r,t)}^{i})).$ (7)
@@ Appendix C Details of applying the trained KGE by MED to real applications
The product recommendation task is to properly recommend items to users that users will interact with a high probability and it often runs on terminal devices. Following PKGM DBLP:conf/icde/ZhangWYWZC21 , which recommends items to users using the neural collaborative filtering (NCF) DBLP:conf/www/HeLZNHC17 framework with the help of pre-trained user embeddings as service vectors, we add trained user embeddings over SKG as service vectors to NCF. In NCF, the MLP layer is used to learn item-user interactions based on the latent feature of the user and item, that is, for a given user-item pair $user_{i}-item_{j}$, the interaction function is
@@
and the other parts of NCF stay the same as in PKGM DBLP:conf/icde/ZhangWYWZC21 .
We train entity and relation embeddings for SKG based on TransE DBLP:conf/nips/BordesUGWY13 and input the trained entity (user) embedding into Equation (9) and Equation (11).
@@ D.1 Dataset and Evaluation Metric
For the experiments extending MED to BERT, we adopt the common GLUE DBLP:conf/iclr/WangSMHLB19 benchmark for evaluation. To be specific, we use the development set of the GLUE benchmark which includes four tasks: Paraphrase Similarity Matching, Sentiment Classification, Natural Language Inference, and Linguistic Acceptability. For Paraphrase Similarity Matching, we use MRPC DBLP:conf/acl-iwp/DolanB05 , QQP and STS-B DBLP:conf/lrec/ConneauK18 for evaluation. For Sentiment Classification, we use SST-2 DBLP:conf/emnlp/SocherPWCMNP13 . For Natural Language Inference, we use MNLI DBLP:conf/naacl/WilliamsNB18 , QNLI DBLP:conf/emnlp/RajpurkarZLL16 , and RTE for evaluation. In terms of evaluation metrics, we follow previous work DBLP:conf/naacl/DevlinCLT19 ; DBLP:conf/emnlp/SunCGL19 . For MRPC and QQP, we report F1 and accuracy. For STS-B, we consider Pearson and Spearman correlation as our metrics. The other tasks use accuracy as the metric. For MNLI, the results of MNLI-m and MNLI-mm are both reported separately.
@@ D.2 Baselines
For comparison, we choose Knowledge Distillation (KD) models and Hardware-Aware Transformers DBLP:conf/acl/WangWLCZGH20 (HAT) customized for transformers as baselines. For the KD models, we compare MED with Basic KD (BKD) DBLP:journals/corr/HintonVD15 , Patient KD (PKD) DBLP:conf/emnlp/SunCGL19 , Relational Knowledge Distillation (RKD) DBLP:conf/cvpr/ParkKLC19 , Deep Self-attention Distillation (MiniLM) DBLP:conf/nips/WangW0B0020 , Meta Learning-based KD (MetaDistill) DBLP:conf/acl/ZhouXM22 and Feature Structure Distillation (FSD) DBLP:journals/eswa/JungKNK23 . For the comparability of the results, we choose 4-layer BERT (BERT4) or 6-layer BERT (BERT6) as the student model architectures, which guarantees that the number of model parameters (#P(M)) or speedup is comparable. For HAT, we use the same model architecture as our MED for training and show the results of sub-models with three parameter scales.
@@ D.3 Implementation
To implement MED on BERT, for the word embedding layer, all sub-models share the front portion of embedding parameters in the same way as in KGE, and for the transformer layer, all sub-models share the front portion of weight parameters as in HAT DBLP:conf/acl/WangWLCZGH20 . Specifically, assuming that the embedding dimension of the largest BERT model $B_{n}$ is $d_{n}$, and the embedding dimension of the sub-model $B_{i}$ is $d_{i}$, for any parameter matrix with the shape $x\times y$ in $B_{n}$, the front portion sub-matrix of it with the shape $\frac{d_{i}}{d_{n}}x\times\frac{d_{i}}{d_{n}}y$ is the parameter matrix of the corresponding position in $B_{i}$. Finally, it just need to replace the triple score $s_{(h,r,t)}$ in Equation (1), Equation (2), Equation (3), and Equation (4) with the logits output for the corresponding category of the classifier in the classification task.

### v2

@@ Appendix A More Results of link prediction
TransE Method MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10
DT 0.150 0.235 0.277 0.440 0.299 0.477 0.313 0.484 0.315 0.499 0.318 0.501 0.322 0.508
Ext 0.115 0.211 0.191 0.324 0.236 0.392 0.266 0.436 0.286 0.462 0.299 0.479 0.322 0.508
Ext-L 0.109 0.194 0.175 0.293 0.232 0.381 0.263 0.424 0.285 0.462 0.301 0.484 0.322 0.508
Ext-V 0.139 0.256 0.200 0.348 0.237 0.396 0.270 0.437 0.293 0.466 0.308 0.488 0.322 0.508
BKD 0.176 0.293 0.279 0.446 0.303 0.480 0.315 0.500 0.315 0.501 0.320 0.502 - -
TA 0.175 0.246 0.281 0.441 0.303 0.484 0.314 0.498 0.319 0.504 0.321 0.504 - -
DualDE 0.179 0.301 0.281 0.443 0.306 0.483 0.316 0.502 0.319 0.505 0.322 0.508 - -
IterDE 0.176 0.285 0.276 0.446 0.307 0.482 0.315 0.503 0.317 0.505 0.319 0.505 - -
MED 0.196 0.341 0.290 0.472 0.308 0.486 0.317 0.502 0.320 0.505 0.321 0.507 0.322 0.507
SimplE Method MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10
DT 0.097 0.179 0.176 0.321 0.236 0.390 0.271 0.431 0.285 0.458 0.291 0.467 0.295 0.472
Ext 0.037 0.068 0.069 0.107 0.090 0.144 0.159 0.258 0.229 0.372 0.269 0.432 0.295 0.472
Ext-L 0.045 0.059 0.056 0.062 0.083 0.146 0.114 0.205 0.196 0.316 0.258 0.421 0.295 0.472
Ext-V 0.049 0.069 0.066 0.101 0.105 0.149 0.138 0.224 0.224 0.369 0.261 0.414 0.295 0.472
BKD 0.113 0.204 0.182 0.315 0.244 0.412 0.275 0.439 0.287 0.463 0.293 0.470 - -
TA 0.124 0.221 0.192 0.329 0.254 0.416 0.276 0.448 0.290 0.465 0.295 0.471 - -
DualDE 0.120 0.213 0.195 0.346 0.258 0.429 0.279 0.443 0.293 0.466 0.296 0.468 - -
IterDE 0.120 0.215 0.193 0.338 0.257 0.427 0.281 0.440 0.293 0.465 0.297 0.468 - -
MED 0.143 0.267 0.233 0.384 0.261 0.427 0.279 0.448 0.291 0.466 0.293 0.468 0.294 0.470
RotatE Method MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10
DT 0.254 0.424 0.297 0.477 0.312 0.495 0.317 0.502 0.322 0.506 0.323 0.510 0.325 0.515
Ext 0.138 0.245 0.203 0.340 0.251 0.410 0.276 0.443 0.291 0.465 0.305 0.485 0.325 0.515
Ext-L 0.135 0.243 0.188 0.319 0.221 0.365 0.246 0.402 0.280 0.453 0.299 0.477 0.325 0.515
Ext-V 0.160 0.281 0.198 0.340 0.238 0.393 0.265 0.427 0.288 0.458 0.302 0.478 0.325 0.515
BKD 0.277 0.442 0.305 0.485 0.314 0.503 0.321 0.508 0.322 0.510 0.323 0.509 - -
TA 0.280 0.447 0.306 0.485 0.313 0.501 0.319 0.507 0.323 0.510 0.323 0.509 - -
DualDE 0.282 0.449 0.307 0.486 0.315 0.502 0.318 0.507 0.322 0.512 0.324 0.514 - -
IterDE 0.276 0.445 0.306 0.482 0.317 0.504 0.319 0.508 0.323 0.512 0.324 0.513 - -
MED 0.288 0.459 0.311 0.492 0.318 0.504 0.322 0.509 0.323 0.510 0.324 0.512 0.324 0.514
PairRE Method MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10
DT 0.182 0.314 0.243 0.395 0.284 0.452 0.307 0.476 0.319 0.505 0.328 0.518 0.332 0.522
Ext 0.148 0.222 0.177 0.289 0.217 0.353 0.259 0.416 0.294 0.469 0.321 0.506 0.332 0.522
Ext-L 0.150 0.249 0.196 0.294 0.219 0.333 0.271 0.436 0.309 0.489 0.326 0.513 0.332 0.522
Ext-V 0.176 0.277 0.192 0.303 0.229 0.374 0.279 0.450 0.311 0.490 0.329 0.513 0.332 0.522
BKD 0.198 0.332 0.251 0.407 0.288 0.453 0.311 0.487 0.321 0.508 0.330 0.521 - -
TA 0.208 0.346 0.263 0.430 0.292 0.455 0.314 0.493 0.323 0.509 0.332 0.521 - -
DualDE 0.207 0.342 0.261 0.427 0.293 0.456 0.316 0.495 0.326 0.512 0.334 0.524 - -
IterDE 0.205 0.340 0.264 0.431 0.293 0.462 0.314 0.494 0.324 0.508 0.332 0.522 - -
MED 0.239 0.384 0.274 0.437 0.303 0.466 0.314 0.495 0.324 0.510 0.329 0.521 0.330 0.520
Table 11: MRR and Hit@10 of some representative dimensions on FB15K237.
@@
TransE Method Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1
DT 0.169 0.102 0.301 0.190 0.327 0.212 0.340 0.218 0.348 0.222 0.353 0.224 0.358 0.228
Ext 0.123 0.065 0.211 0.122 0.264 0.156 0.296 0.180 0.320 0.197 0.331 0.208 0.358 0.228
Ext-L 0.118 0.065 0.192 0.115 0.256 0.157 0.292 0.180 0.316 0.198 0.333 0.210 0.358 0.228
Ext-V 0.150 0.081 0.222 0.126 0.265 0.156 0.301 0.185 0.325 0.205 0.341 0.217 0.358 0.228
BKD 0.178 0.106 0.308 0.198 0.336 0.208 0.349 0.222 0.349 0.223 0.354 0.226 - -
TA 0.188 0.112 0.307 0.200 0.336 0.212 0.348 0.220 0.353 0.225 0.355 0.223 - -
DualDE 0.193 0.115 0.307 0.201 0.337 0.216 0.351 0.223 0.354 0.226 0.356 0.227 - -
IterDE 0.187 0.112 0.299 0.185 0.333 0.214 0.351 0.222 0.353 0.223 0.354 0.224 - -
MED 0.215 0.122 0.321 0.199 0.338 0.218 0.347 0.223 0.351 0.226 0.356 0.227 0.358 0.227
SimplE Method Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1
DT 0.103 0.055 0.193 0.105 0.256 0.161 0.297 0.191 0.314 0.197 0.323 0.208 0.324 0.211
Ext 0.039 0.019 0.071 0.047 0.091 0.057 0.171 0.109 0.251 0.159 0.294 0.187 0.324 0.211
Ext-L 0.043 0.035 0.048 0.037 0.111 0.040 0.131 0.093 0.216 0.134 0.281 0.177 0.324 0.211
Ext-V 0.047 0.036 0.074 0.043 0.097 0.077 0.145 0.109 0.248 0.156 0.289 0.189 0.324 0.211
BKD 0.123 0.064 0.201 0.115 0.261 0.164 0.299 0.191 0.308 0.202 0.318 0.213 - -
TA 0.133 0.073 0.210 0.123 0.276 0.175 0.302 0.195 0.318 0.203 0.323 0.211 - -
DualDE 0.130 0.071 0.224 0.115 0.279 0.175 0.305 0.196 0.324 0.208 0.326 0.211 - -
IterDE 0.132 0.069 0.217 0.118 0.276 0.174 0.303 0.192 0.326 0.204 0.324 0.212 - -
MED 0.164 0.073 0.254 0.158 0.288 0.177 0.305 0.196 0.319 0.205 0.318 0.209 0.322 0.209
RotatE Method Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1
DT 0.284 0.168 0.330 0.207 0.346 0.223 0.352 0.224 0.353 0.229 0.357 0.230 0.363 0.234
Ext 0.152 0.080 0.225 0.129 0.278 0.170 0.304 0.190 0.322 0.203 0.335 0.217 0.363 0.234
Ext-L 0.147 0.078 0.209 0.121 0.247 0.146 0.275 0.166 0.312 0.193 0.333 0.209 0.363 0.234
Ext-V 0.174 0.097 0.218 0.126 0.264 0.159 0.293 0.182 0.319 0.201 0.336 0.213 0.363 0.234
BKD 0.306 0.193 0.338 0.214 0.352 0.224 0.354 0.230 0.356 0.230 0.358 0.231 - -
TA 0.308 0.196 0.339 0.216 0.353 0.225 0.358 0.229 0.359 0.229 0.358 0.231 - -
DualDE 0.311 0.197 0.341 0.216 0.353 0.227 0.360 0.230 0.361 0.232 0.361 0.233 - -
IterDE 0.307 0.195 0.342 0.215 0.355 0.225 0.359 0.232 0.363 0.233 0.362 0.234 - -
MED 0.324 0.201 0.344 0.216 0.355 0.225 0.357 0.231 0.358 0.233 0.362 0.233 0.362 0.232
PairRE Method Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1 Hit@3 Hit@1
DT 0.198 0.116 0.262 0.162 0.312 0.202 0.337 0.222 0.352 0.227 0.364 0.235 0.368 0.237
Ext 0.158 0.107 0.187 0.118 0.236 0.149 0.283 0.182 0.325 0.207 0.354 0.230 0.368 0.237
Ext-L 0.159 0.099 0.196 0.134 0.238 0.159 0.298 0.188 0.342 0.219 0.359 0.233 0.368 0.237
Ext-V 0.181 0.116 0.192 0.125 0.250 0.154 0.307 0.193 0.343 0.221 0.362 0.237 0.368 0.237
BKD 0.215 0.132 0.265 0.168 0.314 0.203 0.343 0.224 0.355 0.233 0.366 0.236 - -
TA 0.226 0.139 0.291 0.182 0.316 0.210 0.347 0.224 0.358 0.232 0.368 0.235 - -
DualDE 0.224 0.139 0.286 0.179 0.318 0.212 0.351 0.226 0.359 0.234 0.371 0.238 - -
IterDE 0.225 0.135 0.293 0.185 0.324 0.212 0.352 0.224 0.357 0.234 0.369 0.236 - -
MED 0.253 0.172 0.299 0.189 0.327 0.213 0.346 0.224 0.357 0.232 0.366 0.236 0.368 0.235
Table 12: Hit@3 and Hit@1 of some representative dimensions on FB15K237.
@@
(b) SimplE on WN18RR
(c) RotatE on WN18RR
(d) PairRE on WN18RR
(e) TransE on FB15K237
(f) SimplE on FB15K237
(g) RotatE on FB15K237
(h) PairRE on FB15K237
@@ B.2 Evolutionary Improvement Mechanism (EIM)
$L_{KGE}^{i}=\sum_{(h,r,t)\in\mathcal{T}\cup\mathcal{T}^{-}}y\log\sigma(s_{(h,r,t)}^{i})$ (8)
$+(1-y)\log(1-\sigma(s_{(h,r,t)}^{i})).$
@@ Appendix C Details of applying the trained KGE by MED to real applications
The product recommendation task is to properly recommend items to users that users will interact with a high probability and it often runs on terminal devices. Following PKGM Zhang et al. (2021), which recommends items to users using the neural collaborative filtering (NCF) He et al. (2017) framework with the help of pre-trained user embeddings as service vectors, we add trained user embeddings over SKG as service vectors to NCF. In NCF, the MLP layer is used to learn item-user interactions based on the latent feature of the user and item, that is, for a given user-item pair $user_{i}-item_{j}$, the interaction function is
@@
and the other parts of NCF stay the same as in PKGM Zhang et al. (2021).
We train entity and relation embeddings for SKG based on TransE Bordes et al. (2013) and input the trained entity (user) embedding into Equation (10) and Equation (12).
Figure 7: Clustering on FB15K237 with RotatE.
@@ D.1 Dataset and Evaluation Metric
For the experiments extending MED to BERT, we adopt the common GLUE Wang et al. (2019) benchmark for evaluation. To be specific, we use the development set of the GLUE benchmark which includes four tasks: Paraphrase Similarity Matching, Sentiment Classification, Natural Language Inference, and Linguistic Acceptability. For Paraphrase Similarity Matching, we use MRPC Dolan and Brockett (2005), QQP and STS-B Conneau and Kiela (2018) for evaluation. For Sentiment Classification, we use SST-2 Socher et al. (2013). For Natural Language Inference, we use MNLI Williams et al. (2018), QNLI Rajpurkar et al. (2016), and RTE for evaluation. In terms of evaluation metrics, we follow previous work Devlin et al. (2019); Sun et al. (2019a). For MRPC and QQP, we report F1 and accuracy. For STS-B, we consider Pearson and Spearman correlation as our metrics. The other tasks use accuracy as the metric. For MNLI, the results of MNLI-m and MNLI-mm are both reported separately.
@@ D.2 Baselines
For comparison, we choose Knowledge Distillation (KD) models and Hardware-Aware Transformers Wang et al. (2020a) (HAT) customized for transformers as baselines. For the KD models, we compare MED with Basic KD (BKD) Hinton et al. (2015), Patient KD (PKD) Sun et al. (2019a), Relational Knowledge Distillation (RKD) Park et al. (2019), Deep Self-attention Distillation (MiniLM) Wang et al. (2020b), Meta Learning-based KD (MetaDistill) Zhou et al. (2022a) and Feature Structure Distillation (FSD) Jung et al. (2023). For the comparability of the results, we choose 4-layer BERT (BERT4) or 6-layer BERT (BERT6) as the student model architectures, which guarantees that the number of model parameters (#P(M)) or speedup is comparable. For HAT, we use the same model architecture as our MED for training and show the results of sub-models with three parameter scales.
@@ D.3 Implementation
To implement MED on BERT, for the word embedding layer, all sub-models share the front portion of embedding parameters in the same way as in KGE, and for the transformer layer, all sub-models share the front portion of weight parameters as in HAT Wang et al. (2020a). Specifically, assuming that the embedding dimension of the largest BERT model $B_{n}$ is $d_{n}$, and the embedding dimension of the sub-model $B_{i}$ is $d_{i}$, for any parameter matrix with the shape $x\times y$ in $B_{n}$, the front portion sub-matrix of it with the shape $\frac{d_{i}}{d_{n}}x\times\frac{d_{i}}{d_{n}}y$ is the parameter matrix of the corresponding position in $B_{i}$. Finally, it just need to replace the triple score $s_{(h,r,t)}$ in Equation (2), Equation (3), Equation (4), and Equation (5) with the logits output for the corresponding category of the classifier in the classification task.
@@
Appendix E Visual analysis of embedding
We select four primary entity categories (‘organization’, ‘sports’, ‘location’, and ‘music’) that contain more than 300 entities in FB15K237, and randomly select 250 entities for each. We cluster these entities’ embeddings of 3 different dimensions ($d$=10, 100, 600) by the t-SNE algorithm, and the clustering results are visualized in Fig. 7. Under the same dimension, the clustering result of MED is always the best, followed by DualDE, while the result of Ext-V is generally poor, which is consistent with the conclusion in Section 5.2. We also find some special phenomenons for MED when dimension increases: 1) the nodes of the ‘sports’ gradually become two clusters meaning MED learns more fine-grained category information as dimension increases, and 2) the relative distribution among different categories hardly changes and shows a trend of “inheritance” and “improvement”. This further proves MED achieves our expectation that high-dimensional sub-models retain the ability of low-dimensional sub-models, and can learn more knowledge than low-dimensional sub-models.